dx/(z^2+2y)=dy/(z^2+2x)=dz/-z


1
Expert's answer
2021-11-02T12:36:43-0400

"\frac{dx-dy}{2y-2x}=\frac{dz}{-z}"


"ln(x-y)\/2=lnz+lnc'_1"


"c_1=\frac{x-y}{z^2}"


"\frac{xdx-ydy}{z^2(x-y)}=\frac{xdx-ydy}{c_1z^4}=\frac{dz}{-z}"


"x^2-y^2=-2c_1z^4\/4=-c_1z^4\/2"


"c_2=-c_1/2=\frac{x^2-y^2}{z^4}"


"F(c_1,c_2)=F(\frac{x-y}{z^2}, \frac{x^2-y^2}{z^4})=0"